OpenStudy (anonymous):

*** I will give you a medal to help me *** Simplify the rational expression. State any restrictions on the variable. (n^4-11n^2+30)/(n^4-7n^2+10)

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

find factors \[\huge \frac{(n^2-6)(n^2-5)}{(n^2-2)(n^2-5)}\] \[\huge \frac{(n^2-6)( \cancel{n^2-5)}}{(n^2-2)(\cancel{n^2-5})}\] \[\huge \frac{n^2-6}{n^2-2}\]

OpenStudy (anonymous):

what are the restrictions ?

OpenStudy (anonymous):

\[n^2 \ne 2 \implies n \ne \sqrt{2}\]

OpenStudy (anonymous):

\[n \ne \pm \sqrt{2}\]